# references.bib

### v2

@@ Abstract
Single-image depth estimation is essential for endoscopy tasks such as localization, reconstruction, and augmented reality. Most existing methods in surgical scenes focus on in-domain depth estimation, limiting their real-world applicability. This constraint stems from the scarcity and inferior labeling quality of medical data for training. In this work, we present EndoOmni, the first foundation model for zero-shot cross-domain depth estimation for endoscopy. To harness the potential of diverse training data, we refine the advanced self-learning paradigm that employs a teacher model to generate pseudo-labels, guiding a student model trained on large-scale labeled and unlabeled data. To address training disturbance caused by inherent noise in depth labels, we propose a robust training framework that leverages both depth labels and estimated confidence from the teacher model to jointly guide the student model training. Moreover, we propose a weighted scale-and-shift invariant loss to adaptively adjust learning weights based on label confidence, thus imposing learning bias towards cleaner label pixels while reducing the influence of highly noisy pixels. Experiments on zero-shot relative depth estimation show that our EndoOmni improves state-of-the-art methods in medical imaging for 41% and existing foundation models for 25% in terms of absolute relative error on specific dataset. Furthermore, our model provides strong initialization for fine-tuning to metric depth estimation, maintaining superior performance in both in-domain and out-of-domain scenarios. The source code will be publicly available.
@@ I Introduction
Addressing the limitation of data scarcity, we have assembled the most extensive meta-dataset for medical image SDE, encompassing over 200,000 labeled frames from public and proprietary datasets and an additional 500,000 unlabeled frames (Table I) for model training and evaluation. To harness the potential of unlabeled data, we leverage an self-learning strategy that utilizes a pretrained teacher model to generate pseudo labels. This combined labeled and pseudo-labeled data is then used to train a student model that processes highly perturbed images. This self-learning framework enhances training data variety, promoting generalization by exposing the model to diverse features and patterns. However, since the teacher model predicts for out-of-domain datasets, the resulting pseudo labels are often noisy, posing an additional challenge for robust training.
@@
In this work, we pursue a refined self-learning framework and a robust training loss to guide network training by noisy labels and pseudo labels. We creatively extend the line of thinking in robust training on classification problems [zhou2020robust, liu2020early, li2023disc] to the challenge of depth estimation. Through our empirical experiments, we observe that loss on clean data converges faster due to the inherent consistency of clean labels. Additionally, pseudo-labels tend to be more reliable when the teacher model predicts consistent results for the same instance across different augmentations. Based on this analysis, we propose estimating per-pixel label confidence according to the teacher model’s learning behavior, and using both labels and confidence to jointly guide the training of student model. Moreover, we have developed a robust SDE training loss that adjusts learning weights based on estimated label confidence. This approach not only mitigates training instability caused by label noise but also maintains the contribution of difficult regions in each instance, thereby enhancing the learning effectiveness of the SDE network.
@@
Dataset Organs Size
@@
Hamlyn Centre Datasets Kidney 91,866 fr.
C3VD [bobrow2023] Colon phantom 8,742 fr.
SERV-CT [edwards2020serv] Liver, kidney, heart 32 fr.
EndoAbS [penza2018endoabs] Liver, kidney, spleen 120 fr.
SCARED [allan2021stereo] Abdomen 22,950 fr.
SimColon [rau2019implicit] Colon simulation 16,016 fr.
EndoMapper [azagra2023endomapper] Colon simulation 1,919 fr.
Ours-Phantom Airway phantom 395 fr.
Ours-Patient Airway 70,223 fr.
@@
EndoSLAM [ozyoruk2021endoslam] Colon, stomach, etc. 42,459 fr.
EndoMapper [azagra2023endomapper] Colon 241,887 fr.
ROBUST-MIS [ross2021comparative] Abdomen 128,501 fr.
EAD [ali2020objective] Stomach, ureter, etc. 2,531 fr.
Surgical-Vis [zia2021surgical] - 3,360 fr.
CholecT50 [nwoye2022rendezvous] Abdomen 100,863 fr.
CVC-ClinicDB [bernal2015wm] Colon 612 fr.
  • Dataset size refers to the number of image frames actually used. Original datasets might be selectively extracted to ensure variability.
@@ II Related Works
SDE in Medical Endoscopy. SDE in endoscopic images aids in precise navigation, accurate lesion assessment, and improved surgical planning. Given the impracticality of using depth cameras or LiDAR in such settings, most datasets depend on correlating endoscopic views with pre-operative CT scans or generating sparse point clouds via calibrated stereo-endoscopes. However, the limited availability of training data remains a significant barrier. As a result, existing research of SDE in medical imaging put extra effort into learning through view synthesis [recasens2021, shao2022self, liu2019dense, ozyoruk2021endoslam] and generative models [tian2024dd, mathew2020augmenting, shen2019context, wei2024absolute] to estimate depth for individual datasets or to enhance unsupervised training efficacy. With the success of foundation models and their outstanding performance in zero-shot SDE, a few recent works attempted to adapt foundation models trained on natural images to medical imaging [cui2024surgical, cui2024endodac]. However, due to the large domain gap between medical images and natural images, zero-shot generalization of these foundation models are not as good as expected [han2024depth].
@@
In this study, we have curated the largest meta-dataset for endoscopy SDE and trained our EndoOmni model on this extensive mixed dataset. To mitigate label noise, we introduce a self-learning framework that uses both (pseudo) labels and their associated confidence map to guide network training. Additionally, we propose a weighted scale-and-shift-invariant loss to concentrate gradient updates on regions with high confidence labels.
@@ III-A Teacher Model Training
Several endoscopic datasets with depth annotations [allan2021stereo, bobrow2023, edwards2020serv, azagra2023endomapper, penza2018endoabs, rau2019implicit] are employed. Accurately annotating depth labels in endoscopic data presents a significant challenge compared to general open scenes because high-precision sensors are impractical to set up during clinical procedures. Consequently, existing datasets generate annotations using stereo cameras [allan2021stereo, edwards2020serv, penza2018endoabs], registration to CT scans [bobrow2023], or synthetic data [azagra2023endomapper, rau2019implicit]. To expand the training data, we have curated an additional dataset for bronchoscopy. Depth annotations are sourced using registration-based methods, airway meshes from patient CT scans are reconstructed, and video frames are manually registered to these in a virtual endoscopic view. Depth labels are generated by rendering depth maps in the airway model with the registered endoscopic poses. As a result, our teacher model leverages these extensive datasets, further enhancing training data diversity by generating pseudo-labels for unlabeled datasets to guide the student model training.
@@ III-B Robust Student Model Training
To extend training data diversity for student model, besides labeled data from sources [allan2021stereo, bobrow2023, edwards2020serv, penza2018endoabs, rau2019implicit, azagra2023endomapper], we have incorporated unlabeled endoscopic data from datasets intended for other tasks [ozyoruk2021endoslam, azagra2023endomapper, nwoye2022rendezvous, bernal2015wm, zia2021surgical, ross2021comparative, ali2020objective]. Each dataset introduces unique challenges due to variations in annotation accuracy—depth labels from stereo cameras can be sparse and imprecise due to calibration errors and environmental factors, while labels from registration methods may suffer from errors in registration and CT reconstruction inaccuracies. Leveraging this mix of unlabeled data enhances the breadth of our training data, yet introduces the risk of noisy pseudo-labels from the teacher model, which can mislead the student model’s training. This is particularly problematic in medical imaging, where fewer labeled data are available to train the teacher model, leading to its unstable performance on out-of-distribution unlabeled data. Such noise in training could result in degraded generalization performance during practical application. To address this, we propose to mitigate the impact of these inaccuracies by estimating the confidence of each labeled pixel. We then adjust the learning weight for each pixel based on its estimated confidence. This approach focuses gradient updates on regions with higher confidence, while reducing the influence of gradients from noisier regions.
@@
where the confidence of depth labels are determined by the SSI discrepancy between manual labels and pretrained teacher model prediction based on Eq. 2.
@@
EDM* [recasens2021] 0.185 5.424 16.100 0.255 0.732
AF-SfMLearner* [shao2022self] 0.168 4.44 13.87 0.204 0.77
Surgical-DINO* [cui2024surgical] 0.146 3.216 11.974 0.178 0.801
EndoDAC* [cui2024endodac] 0.138 2.796 11.491 0.171 0.813
Depth-Anything* [yang2024depth] 0.174 3.829 13.226 0.210 0.750
Depth-Anything2* [yang2024depth2] 0.229 5.949 16.718 0.269 0.649
Ours* 0.135 2.467 10.277 0.167 0.826
Ours 0.097 1.286 7.687 0.123 0.875
  • Best and second best performances are highlighted. * denotes recovering scale before evaluation. $\triangle$ denotes recovering scale and shift before evaluation. $\ddagger$ denotes in-domain evaluation.
@@
SfMLearner* [zhou2017unsupervised] 0.151 3.917 17.451 0.191 0.779
Fang et al.* [fang2020towards] 0.149 3.099 15.564 0.188 0.787
DeFeat-Net* [spencer2020defeat] 0.114 1.946 12.588 0.153 0.873
SC-SfMLearner* [bian2019unsupervised] 0.117 2.015 12.415 0.148 0.852
Monodepth2* [godard2019digging] 0.123 2.205 12.927 0.152 0.856
Endo-SfM* [ozyoruk2021endoslam] 0.116 2.014 12.493 0.143 0.864
AF-SfMLearner* [shao2022self] 0.102 1.632 11.092 0.131 0.898
Depth-Anything* [yang2024depth] 0.080 0.906 7.212 0.098 0.928
Depth-Anything2* [yang2024depth2] 0.130 2.125 10.578 0.160 0.818
Ours* 0.060 0.445 5.511 0.075 0.988
Ours 0.034 0.158 3.316 0.044 1.000
  • Best and second best are highlighted. * denotes recovering scale before evaluation. $\triangle$ denotes recovering scale and shift before evaluation.
@@ IV-A Implementation Details
We adopt the DINOv2 [oquab2024dinov2] encoder and DPT [ranftl2021vision] decoder for both the teacher and student models. We first train the teacher model on all labeled training sets until convergence. We then train the student model with a batch size of 32 until convergence under the teacher model’s guidance. Following the approach in [yang2024depth], we challenge the student model with strong perturbations including color jittering and Gaussian blurring. Additionally, a semantic perception technique aligning backbone features with DINOv2 [oquab2024dinov2] is adopted. Labeled data and unlabeled data are mixed in a 1:3 ratio for each input batch. For both teacher and student network training, we initialize the encoder with pretrained weights [yang2024depth] for faster convergence and randomly initialize the decoder weights. We use a learning rate of $5\times 10^{-6}$ and a learning rate factor of 0.1:1 for the encoder and decoder. The networks are trained with the Adam optimizer with a linear decay learning rate.
For a fair comparison with recent work in medical image depth estimation, we fine-tune our foundation model $f$ to produce a refined model $\dot{f}$ for zero-shot relative depth estimation (RDE), leaving only an unknown scale factor. Note that the key contribution remains the foundation model $f$. We fine-tune the student model $f$ with a frozen encoder on the SCARED dataset using scale-invariant loss. SCARED was selected for its high annotation accuracy and sufficient training frames. Formally, the scale-invariant loss[nips2014_7bccfde7] is defined as:
@@
During experiments, we included our scale-invariant fine-tuned version of model $\dot{f}$ using the same median scaling preprocessing as existing methods to ensure a fair comparison.
@@
Ours 0.345 0.987 2.568 0.337 0.519
Figure 5: Zero-shot performance of EndoOmni on SERV-CT (top two rows) and the Hamlyn Dataset (bottom two rows), compared with EndoDAC [cui2024endodac], the leading SOTA method for endoscopy, and Depth Anything [yang2024depth], the top-performing SOTA foundation model. Quantitative results are also provided for the model without our robust training loss, denoted as SSI. EndoOmni demonstrates the most accurate geometric structure and the sharpest edges among all methods.
@@
Depth-Anything [yang2024depth] 0.508 4.280 9.430 0.576 0.329
Ours 0.364 1.914 5.607 0.411 0.382
@@ IV-B Evaluation Metrics
We assess our approach using five widely recognized metrics in depth estimation: AbsRel (absolute relative error), SqRel (square relative error), RMSE (root mean square error), RMSElog (root mean square logarithmic error), and $\delta_{1}$, following established works [ranftl2020towards, yang2024depth, recasens2021, cui2024endodac].
@@
For RDE comparison with existing medical imaging research, we perform median scaling [zhou2017unsupervised] for the fine-tuned version model $\dot{f}$ expressed by:

### v3

@@ Abstract
Single-image depth estimation is essential for endoscopy tasks such as localization, reconstruction, and augmented reality. Most existing methods in surgical scenes focus on in-domain depth estimation, limiting their real-world applicability. This constraint stems from the scarcity and inferior labeling quality of medical data for training. In this work, we present EndoOmni, the first foundation model for zero-shot cross-domain depth estimation for endoscopy. To harness the potential of diverse training data, we refine the advanced self-learning paradigm that employs a teacher model to generate pseudo-labels, guiding a student model trained on large-scale labeled and unlabeled data. To address training disturbance caused by inherent noise in depth labels, we propose a robust training framework that leverages both depth labels and estimated confidence from the teacher model to jointly guide the student model training. Moreover, we propose a weighted scale-and-shift invariant loss to adaptively adjust learning weights based on label confidence, thus imposing learning bias towards cleaner label pixels while reducing the influence of highly noisy pixels. Experiments on zero-shot relative depth estimation show that our EndoOmni improves state-of-the-art methods in medical imaging for 33% and existing foundation models for 34% in terms of absolute relative error on specific datasets. Furthermore, our model provides strong initialization for fine-tuning metric depth estimation, maintaining superior performance in both in-domain and out-of-domain scenarios. The source code is publicly available at https://github.com/TianCuteQY/EndoOmni.
@@ I Introduction
Addressing the limitation of data scarcity, we have assembled the most extensive meta-dataset for medical image SDE, encompassing over 200,000 labeled frames from public and proprietary datasets and an additional 500,000 unlabeled frames (Table I) for model training and evaluation. To harness the potential of unlabeled data, we leverage a self-learning strategy that utilizes a pretrained teacher model to generate pseudo labels. This combined labeled and pseudo-labeled data is then used to train a student model that processes highly perturbed images. This self-learning framework enhances training data variety, promoting generalization by exposing the model to diverse features and patterns. However, since the teacher model predicts out-of-domain datasets, the resulting pseudo labels are often noisy, posing an additional challenge for robust training.
@@
In this work, we pursue a refined self-learning framework and a robust training loss to guide network training by noisy labels and pseudo labels. We creatively extend the line of thinking in robust training on classification problems [zhou2020robust, liu2020early, li2023disc] to the challenge of depth estimation. Through our empirical experiments, we observe that loss on clean data converges faster due to the inherent consistency of clean labels. Additionally, pseudo-labels tend to be more reliable when the teacher model predicts consistent results for the same instance across different augmentations. Based on this analysis, we propose estimating per-pixel label confidence according to the teacher model’s learning behavior, and using both labels and confidence to jointly guide the training of the student model. Moreover, we have developed a robust SDE training loss that adjusts learning weights based on estimated label confidence. This approach not only mitigates training instability caused by label noise but also maintains the contribution of difficult regions in each instance, thereby enhancing the learning effectiveness of the SDE network.
@@
Dataset Organs Size Usage
@@
Hamlyn Centre Datasets Kidney 91,866 fr. E
C3VD [bobrow2023] Colon phantom 10,015 fr. T
SERV-CT [edwards2020serv] Liver, kidney, heart 32 fr. E
EndoAbS [penza2018endoabs] Liver, kidney, spleen 240 fr. T
SCARED [allan2021stereo] Abdomen 17,056 fr. T
SimColon [rau2019implicit] Colon simulation 16,016 fr. T
EndoMapper [azagra2023endomapper] Colon simulation 1,919 fr. T
Ours-Phantom Airway phantom 395 fr. E
Ours-Patient Airway 70,223/5,754 fr. T/E
@@
EndoSLAM [ozyoruk2021endoslam] Colon, stomach, etc. 42,459 fr. T
EndoMapper [azagra2023endomapper] Colon 241,887 fr. T
ROBUST-MIS [ross2021comparative] Abdomen 128,501 fr. T
EAD [ali2020objective] Stomach, ureter, etc. 2,531 fr. T
Surgical-Vis [zia2021surgical] - 3,360 fr. T
CholecT50 [nwoye2022rendezvous] Abdomen 100,863 fr. T
CVC-ClinicDB [bernal2015wm] Colon 612 fr. T
  • Dataset size refers to the number of image frames actually used for training or testing. Original datasets may be selectively extracted to ensure variability. T and E refer to the train and evaluation datasets, respectively. For Ours-Patient, only the training splits are used for training the foundation model, excluding the test splits.
@@ II Related Works
SDE in Medical Endoscopy. SDE in endoscopic images aids in precise navigation, accurate lesion assessment, and improved surgical planning. Given the impracticality of using depth cameras or LiDAR in such settings, most datasets depend on correlating endoscopic views with pre-operative CT scans or generating sparse point clouds via calibrated stereo-endoscopes. However, the limited availability of training data remains a significant barrier. As a result, existing research of SDE in medical imaging put extra effort into learning through view synthesis [recasens2021, shao2022self, liu2019dense, ozyoruk2021endoslam] and generative models [tian2024dd, mathew2020augmenting, shen2019context, wei2024absolute] to estimate depth for individual datasets or to enhance unsupervised training efficacy. With the success of foundation models and their outstanding performance in zero-shot SDE, a few recent works attempted to adapt foundation models trained on natural images to medical imaging [cui2024surgical, cui2024endodac]. However, due to the large domain gap between medical images and natural images, zero-shot generalization of these foundation models is not as good as expected [han2024depth].
@@
In this study, we have curated the largest meta-dataset for endoscopy SDE and trained our EndoOmni model on this extensive mixed dataset. To mitigate label noise, we introduce a self-learning framework that uses both (pseudo) labels and their associated confidence map to guide network training. Additionally, we propose a weighted scale-and-shift-invariant loss to concentrate gradient updates on regions with high-confidence labels.
@@ III-A Teacher Model Training
The teacher model is trained with all labeled data listed in Table I with notation ”T”. Accurately annotating depth labels in endoscopic data presents a significant challenge compared to general open scenes because high-precision sensors are impractical to set up during clinical procedures. Consequently, existing datasets generate annotations using stereo cameras [allan2021stereo, edwards2020serv, penza2018endoabs], registration to CT scans [bobrow2023], or synthetic data [azagra2023endomapper, rau2019implicit]. To expand the training data, we have curated an additional dataset for bronchoscopy. Depth annotations are sourced using registration-based methods, airway meshes from patient CT scans are reconstructed, and video frames are manually registered to these in a virtual endoscopic view. Depth labels are generated by rendering depth maps in the airway model with the registered endoscopic poses. As a result, our teacher model leverages these extensive datasets, further enhancing training data diversity by generating pseudo-labels for unlabeled datasets to guide the student model training.
@@ III-B Robust Student Model Training
To extend training data diversity for the student model, besides labeled data, we have incorporated unlabeled endoscopic data from datasets intended for other tasks. All datasets used for training and evaluation of the student model are listed in Table I with notation ”T” and ”E”. Each dataset introduces unique challenges due to variations in annotation accuracy—depth labels from stereo cameras can be sparse and imprecise due to calibration errors and environmental factors, while labels from registration methods may suffer from errors in registration and CT reconstruction inaccuracies. Leveraging this mix of unlabeled data enhances the breadth of our training data, yet introduces the risk of noisy pseudo-labels from the teacher model, which can mislead the student model’s training. This is particularly problematic in medical imaging, where fewer labeled data are available to train the teacher model, leading to its unstable performance on out-of-distribution unlabeled data. Such noise in training could result in degraded generalization performance during practical application. To address this, we propose to mitigate the impact of these inaccuracies by estimating the confidence of each labeled pixel. We then adjust the learning weight for each pixel based on its estimated confidence. This approach focuses gradient updates on regions with higher confidence, while reducing the influence of gradients from noisier regions.
@@
where the confidence of depth labels is determined by the SSI discrepancy between manual labels and pretrained teacher model prediction based on Eq. 2.
@@
EDM [recasens2021] 0.185 5.424 16.100 0.255 0.732
AF-SfMLearner [shao2022self] 0.168 4.440 13.870 0.204 0.770
Surgical-DINO [cui2024surgical] 0.146 3.216 11.974 0.178 0.801
EndoDAC [cui2024endodac] 0.138 2.796 11.491 0.171 0.813
IID-SfmLearner [li2024image] 0.157 2.995 11.621 0.192 0.776
DepthAnything-L [yang2024depth] 0.174 3.829 13.226 0.210 0.750
DepthAnything2-L [yang2024depth2] 0.229 5.949 16.718 0.269 0.649
EndoOmni (Ours) 0.125 2.337 10.858 0.160 0.827
  • Best performances are highlighted. All results obtained by recovering scale before evaluation. $\ddagger$ denotes in-domain evaluation.
@@
SfMLearner [zhou2017unsupervised] 0.151 3.917 17.451 0.191 0.779
Fang et al. [fang2020towards] 0.149 3.099 15.564 0.188 0.787
DeFeat-Net [spencer2020defeat] 0.114 1.946 12.588 0.153 0.873
SC-SfMLearner [bian2019unsupervised] 0.117 2.015 12.415 0.148 0.852
Monodepth2 [godard2019digging] 0.123 2.205 12.927 0.152 0.856
Endo-SfM [ozyoruk2021endoslam] 0.116 2.014 12.493 0.143 0.864
AF-SfMLearner [shao2022self] 0.102 1.632 11.092 0.131 0.898
IID-SfmLearner[li2024image] 0.123 1.870 10.985 0.153 0.833
MonoPCC [wang2024monopcc] 0.091 1.252 10.059 0.116 0.915
EndoDAC[cui2024endodac] 0.079 0.983 8.723 0.103 0.945
DepthAnything-L [yang2024depth] 0.080 0.906 7.212 0.098 0.928
DepthAnything2-L [yang2024depth2] 0.130 2.125 10.578 0.160 0.818
EndoOmni (Ours) 0.053 0.432 5.922 0.070 0.988
  • Best performances are highlighted. All results were obtained by recovering the scale before evaluation.
@@ IV-A Implementation Details
We adopt the DINOv2 [oquab2024dinov2] encoder and DPT [ranftl2021vision] decoder for both the teacher and student models, as introduced by [yang2024depth]. We first train the teacher model with a ViT-L backbone [dosovitskiy2021an] on all labeled training sets until convergence. We then train the student model with a batch size of 32 until convergence under the teacher model’s guidance. We offer three student model scales based on ViT-S, ViT-B, and ViT-L architectures respectively. Following the approach in [yang2024depth], we challenge the student model with strong perturbations including color jittering and Gaussian blurring. Additionally, a semantic perception technique aligning backbone features with DINOv2 [oquab2024dinov2] is adopted. Labeled data and unlabeled data are mixed in a 1:3 ratio for each input batch. For both teacher and student network training, we initialize the encoder with pretrained weights [yang2024depth] for faster convergence and randomly initialize the decoder weights. The choice of pretrained Depth Anything weights over alternatives like Depth Anything v2 [yang2024depth2] is based on preliminary experiments. We use a learning rate of $5\times 10^{-6}$ and a learning rate factor of 0.1:1 for the encoder and decoder. The networks are trained with the Adam optimizer with a linear decay learning rate.
For a fair comparison with recent work in medical image depth estimation, we fine-tune our foundation model $f$ to produce a refined model $\hat{f}$ for zero-shot relative depth estimation (RDE), leaving only an unknown scale factor. Note that the key contribution remains the foundation model $f$. The student model $f$ traversed the SCARED dataset once using scale-invariant loss. SCARED was selected for its high annotation accuracy and sufficient training frames. Formally, the scale-invariant loss[nips2014_7bccfde7] is defined as:
@@
During experiments, we included our scale-invariant version of model $\hat{f}$ with ViT-B backbone, which brings the parameter count closer to that of previous methods. We used the same median scaling preprocessing as existing methods on SDE in medical imaging, to ensure a fair comparison.
@@
EndoOmni (Ours) 0.345 0.987 2.568 0.337 0.519
Figure 5: Zero-shot performance of EndoOmni on SERV-CT (top two rows) and the Hamlyn Dataset (bottom two rows), compared with EndoDAC [cui2024endodac], the leading SOTA method for endoscopy, and Depth Anything [yang2024depth], the top-performing SOTA foundation model. Quantitative results are also provided for the model without our robust training loss, denoted as SSI. We show corresponding point clouds rendered from the same viewpoint. Misalignments with ground truth are highlighted in boxes on the bottom row for easy identification. EndoOmni demonstrates the most accurate geometric structure and the sharpest edges among all methods.
@@
Depth Anything [yang2024depth] 0.508 4.280 9.430 0.576 0.329
EndoOmni (Ours) 0.364 1.914 5.607 0.411 0.382
@@ IV-B Evaluation Metrics
We assess our approach using five widely recognized metrics of depth estimation: AbsRel (absolute relative error), SqRel (square relative error), RMSE (root mean square error), RMSElog (root mean square logarithmic error), and $\delta_{1}$, following established works [ranftl2020towards, yang2024depth, recasens2021, cui2024endodac].
@@
For RDE comparison with existing medical imaging research, we perform median scaling [zhou2017unsupervised] for the fine-tuned version model $\hat{f}$ expressed by: